OpenStudy (anonymous):

Can someone help me find out the function of this chart? In slope intercept form please

OpenStudy (anonymous):



OpenStudy (welshfella):

the slope will be difference in value of f(x) / corresponding difference in x slope m = (5-1 )/ (2-1)

OpenStudy (welshfella):

confirm this with (1 - (-7)) / (1 - (-1))

OpenStudy (welshfella):

the general form is y = mx + b where b is the y-intercept plug in a value of x and y ( f(x)) and m to find b

OpenStudy (anonymous):

the diffrences in f(x) is different every point though? The slope is supposed to be a stagnant rate of change

OpenStudy (welshfella):

work those 2 expressions out - you'll see they are the same

OpenStudy (anonymous):

OHhhhhhh I see

OpenStudy (anonymous):

wait maybe not im still confused :(

OpenStudy (welshfella):

(5-1)/(2-1) = 4/1 = 4

OpenStudy (anonymous):

Oh ok now I get it thanks so much dude, something so simple yet I was racking my brain to the point of explosion with it haha

OpenStudy (welshfella):

lol yw